# Invariance Matters: Empowering Social Recommendation via Graph Invariant Learning

### v1

@@ Abstract.
Graph-based social recommendation systems have shown significant promise in enhancing recommendation performance, particularly in addressing the issue of data sparsity in user behaviors. Typically, these systems leverage Graph Neural Networks (GNNs) to capture user preferences by incorporating high-order social influences from observed social networks. However, existing graph-based social recommendations often overlook the fact that social networks are inherently noisy, containing task-irrelevant relationships that can hinder accurate user preference learning. The removal of these redundant social relations is crucial, yet it remains challenging due to the lack of ground truth. In this paper, we approach the social denoising problem from the perspective of graph invariant learning and propose a novel method, Social Graph Invariant Learning(SGIL). Specifically, SGIL aims to uncover stable user preferences within the input social graph, thereby enhancing the robustness of graph-based social recommendation systems. To achieve this goal, SGIL first simulates multiple noisy social environments through graph generators. It then seeks to learn environment-invariant user preferences by minimizing invariant risk across these environments. To further promote diversity in the generated social environments, we employ an adversarial training strategy to simulate more potential social noisy distributions. Extensive experimental results demonstrate the effectiveness of the proposed SGIL. The code is available at https://github.com/yimutianyang/SIGIR2025-SGIL.
@@
copyright: acmlicensedccs: Human-centered computing Collaborative and social computing
@@ 1. Introduction
In this work, we introduce a novel method, Social Graph Invariant Learning(SGIL), to address these challenges. First, we design preference-guided environment generators that simulate multiple noisy environments, thereby approximating real-world scenarios where social noise is prevalent. Next, we capture invariant user preferences by leveraging an invariance-based regularizer. Specifically, we derive the optimization objective for each environment, taking into account the characteristics of the recommendation task, and incorporate an invariance penalty term to enforce the stability of user preferences across environments. To further ensure the diversity of the simulated noisy environments, we employ an adversarial training strategy. This strategy challenges the model iteratively, forcing it to perform well in progressively more difficult and varied environments, which ultimately improves its generalization ability. The main contributions of this work are summarized as follows:
@@
  • We conduct extensive experiments on three benchmark datasets. Empirical studies, including both real and semi-synthetic datasets, demonstrate the effectiveness of the proposed SGIL
    method in improving social recommendation performance and robustness.
@@ 3.2. Invariant Preference Learning
After obtaining the learned node representations in the $k^{th}$ social environment, we can infer the propensity that user $a$ will interacts with item $i$ by an inner product: $\hat{r}_{ai}=<p_{a}^{k},q_{i}^{k}>$. All the above process are summarized as $\hat{r}_{ai}^{k}=f_{\phi}^{k}(a,i,\mathcal{G}_{r},\mathcal{G}_{s}^{k})$.
@@ 4. Experiments
In this section, we conduct extensive experiments on three benchmarks to demonstrate the effectiveness of our proposed SGIL . In particular, we aim to investigate the following Research Questions (RQs):
  • RQ1: How does the proposed SGIL model performs compared with SOTAs?
  • RQ2: How does each component of SGIL affects for social recommendation tasks?
  • RQ3: How does the proposed SGIL model perform in terms of robustness to noises, sparse scenarios, and various parameters?
@@ 4.1.1. Datasets
The evaluation metrics are Recall@N and NDCG@N, which have been widely used in ranking-based recommender systems. Following the mainstream evaluation process (Zhao et al., 2020), we adopt a full-ranking protocol that all non-interacted items are candidates for evaluation.
@@ 4.1.3. Implementation Details
All embedding-based methods are initialized by a normal initializer with a mean value of 0 and a standard variance of 0.01, the embedding size is fixed to 64. For model optimization, we use the Adam optimizer with a learning rate of 0.001 and a batch size of 2048. We empirically set the GCN layer $L=3$ and the temperature coefficient $\tau=0.2$ for all datasets. For the edge observation bias, we follow the previous work (Yang et al., 2024) and set the bias $\epsilon=0.5$ for all datasets. For the environment number $K$, the variance-based penalty coefficient $\beta$, we carefully search the best parameters for each dataset and report the detailed comparisons. For the adversarial training step, we set $T=20$ on both the Douban-Book and Yelp datasets, and $T=3$ on the Epinions datasets. For all baselines, we refer to the original parameters and carefully fine-tune them for fair comparisons.
@@
4.2. Performance Comparisons (RQ1)
To verify the effectiveness of our proposed SGIL, we compare SGIL with competing methods. As shown in Table LABEL:tab:_douban_results-Table LABEL:tab:_epinions_results, we present recommendation performances on three benchmarks. Following the current SOTA method (Yang et al., 2024), we implement all social denoising methods on the LightGCN-S backbone. Given the empirical studies, we have the following observations:
@@ 4.3. Ablation Study (RQ2)
4.4. Detailed Analysis of SGIL (RQ3)
@@
5. Related Works
5.1. Graph-based Social Recommendation
Social recommendations have been widely deployed in online services, which leverage social networks as auxiliary information to capture users’ behavior patterns (Marsden and Friedkin, 1993; Tang et al., 2013). Following the social homophily (McPherson et al., 2001) and social influence theory (Marsden and Friedkin, 1993), social recommendations are devoted to characterizing social relation effects on user preferences, including trust-based preference learning (Guo et al., 2015), social regularization (Jamali and Ester, 2010; Ma et al., 2011), and social influence modeling (Wu et al., 2019). Early works mainly exploit shallow social relations to enhance recommendation, ignoring the complex high-order social structure influences. Users’ interactions can be naturally formulated in graph forms (He et al., 2020; Yang et al., 2023a), many graph-based social recommendations have been proposed (Wu et al., 2019; Yu et al., 2021a; Yang et al., 2023b). Particularly, borrowing the power of representation ability of GNNs, graph-based social recommendations achieved an impressive process. The core premise of this work is that incorporating high-order social neighbors and interacted items into user preference learning enriches the available information, thereby addressing the issue of data sparsity. Current studies mainly focus on modeling social structure in Euclidean space (Wu et al., 2020; Yu et al., 2021a), few attempts also explore the potential of Hyperbolic space learning (Wang et al., 2021b; Yang et al., 2023b). However, despite the effectiveness of social graph modeling, current research often overlooks the challenge of social noise. Despite the effectiveness of social graph modeling, current studies rarely notice the social noise problem. Directly using the observed social graph may lead to sub-optimal results. In this work, we propose a novel social denoising method to alleviate the effect of task-irrelevant social structures in recommendations.
5.2. Recommendation Denoising
5.3. Invariant Learning and Applications
Invariant learning has emerged as a promising approach to address the challenges posed by Out-Of-Distribution (OOD) generalization in machine learning (Arjovsky et al., 2019; Krueger et al., 2021; Creager et al., 2021). The core idea behind invariant learning is to uncover stable feature representations that remain robust under varying environmental conditions, thereby enhancing the reliability of label prediction models. This framework is based on the assumption that the data contains both spurious and invariant features, where spurious features vary unpredictably across different environments, while invariant features remain stable and consistent. The primary objective of invariant learning methods is to optimize models using Invariant Risk Minimization (IRM) by introducing an invariance-based penalty during optimization, which discourages reliance on spurious features and encourages the model to learn stable, generalizable patterns. Invariant learning has been successfully applied across a wide range of machine learning tasks, including image classification (Arjovsky et al., 2019; Creager et al., 2021), graph learning (Li et al., 2022; Chen et al., 2022; Wu et al., 2022a), and recommender systems (Du et al., 2022; Wang et al., 2022; Zhang et al., 2023; Wang et al., 2024; Bai et al., 2024). Especially for recommender systems,  (Du et al., 2022; Bai et al., 2024) learn invariant multimedia representation to improve the robustness of the general and cold-start recommendation scenarios. Besides,  (Wang et al., 2022; Zhang et al., 2023) leverage the invariant learning to address popularity debiasing, and  (Wang et al., 2024) captures the invariant KG subgraph to enhance recommendations. In this work, we introduce the invariant learning principle to the social recommendation, intending to capture stable user preferences that persist across diverse social environments.

### v2

@@ Abstract.
Graph-based social recommender systems have demonstrated great potential in alleviating data sparsity by leveraging high-order user influence embedded in social networks. However, most existing methods rely heavily on the observed social graph, which is often noisy and includes spurious or task-irrelevant connections that can mislead user preference learning. Identifying and removing these noisy relations is crucial but challenging due to the lack of ground-truth annotations. In this paper, we approach the social denoising problem from the perspective of graph invariant learning and propose a novel approach, Social Graph Invariant Learning(SGIL). Specifically, SGIL aims to uncover stable user preferences within the input social graph, thereby enhancing the robustness of graph-based social recommendation systems. To achieve this goal, SGIL first simulates multiple noisy social environments through graph generators. It then seeks to learn environment-invariant user preferences by minimizing invariant risk across these environments. To further promote diversity in the generated social environments, we employ an adversarial training strategy to simulate more potential social noisy distributions. Extensive experimental results demonstrate the effectiveness of the proposed SGIL. The code is available at https://github.com/yimutianyang/SIGIR2025-SGIL.
@@
copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730013isbn: 979-8-4007-1592-1/2025/07ccs: Human-centered computing Collaborative and social computing
@@ 1. Introduction
In this work, we propose a novel Social Graph Invariant Learning(SGIL) , to tackle the above challenges. First, we design preference-guided environment generators that simulate multiple noisy environments, thereby approximating real-world scenarios where social noise is prevalent. Next, we capture invariant user preferences by leveraging an invariance-based regularizer. Specifically, we derive the optimization objective for each environment, taking into account the characteristics of the recommendation task, and incorporate an invariance penalty term to enforce the stability of user preferences across environments. To further ensure the diversity of the simulated noisy environments, we employ an adversarial training strategy. This strategy challenges the model iteratively, forcing it to perform well in progressively more difficult and varied environments, which ultimately improves its generalization ability. The main contributions of this work are summarized as follows:
@@
  • We conduct extensive experiments on three benchmark datasets. Empirical studies, including both real and semi-synthetic datasets, demonstrate the effectiveness of the proposed SGIL method in improving social recommendation performance and robustness.
2. Related Works
2.1. Graph-based Social Recommendation
Social recommendations have been widely deployed in online services, which leverage social networks as auxiliary information to capture users’ behavior patterns (Marsden and Friedkin, 1993; Tang et al., 2013). Following the social homophily (McPherson et al., 2001) and social influence theory (Marsden and Friedkin, 1993), social recommendations are devoted to characterizing social relation effects on user preferences, including trust-based preference learning (Guo et al., 2015), social regularization (Jamali and Ester, 2010; Ma et al., 2011), and social influence modeling (Wu et al., 2019). Early works mainly exploit shallow social relations to enhance recommendation, ignoring the complex high-order social structure influences.
Users’ interactions can be naturally formulated in graph forms (He et al., 2020; Yang et al., 2023a), many graph-based social recommendations have been proposed (Wu et al., 2019; Yu et al., 2021a; Yang et al., 2023b). Particularly, graph-based social recommendations achieved an impressive process by borrowing the power of representation ability of GNNs.The core premise of this work is that incorporating high-order social neighbors and interacted items into user preference learning enriches the available information, thereby addressing the issue of data sparsity. Current studies mainly focus on modeling social structure in Euclidean space (Wu et al., 2020; Yu et al., 2021a), few attempts also explore the potential of Hyperbolic space learning (Wang et al., 2021b; Yang et al., 2023b). However, despite the effectiveness of social graph modeling, current research often overlooks the challenge of social noise. Despite the effectiveness of social graph modeling, current studies rarely notice the social noise problem. Directly using the observed social graph may lead to sub-optimal results. In this work, we propose a novel social denoising method to alleviate the effect of task-irrelevant social structures in recommendations.
2.2. Recommendation Denoising
2.3. Invariant Learning and Applications
Invariant learning has emerged as a promising approach to address the challenges posed by Out-Of-Distribution (OOD) generalization in machine learning (Arjovsky et al., 2019; Krueger et al., 2021; Creager et al., 2021; Yue et al., 2025). The core idea behind invariant learning is to uncover stable feature representations that remain robust under varying environmental conditions, thereby enhancing the reliability of label prediction models. This framework is based on the assumption that the data contains both spurious and invariant features, where spurious features vary unpredictably across different environments, while invariant features remain stable and consistent. The primary objective of invariant learning methods is to optimize models using Invariant Risk Minimization (IRM) by introducing an invariance-based penalty during optimization, which discourages reliance on spurious features and encourages the model to learn stable, generalizable patterns. Invariant learning has been successfully applied across a wide range of machine learning tasks, including image classification (Arjovsky et al., 2019; Creager et al., 2021), graph learning (Li et al., 2022; Chen et al., 2022; Wu et al., 2022a), and recommender systems (Du et al., 2022; Wang et al., 2022; Zhang et al., 2023; Wang et al., 2024; Bai et al., 2024). Especially for recommender systems,  (Du et al., 2022; Bai et al., 2024) learn invariant multimedia representation to improve the robustness of the general and cold-start recommendation scenarios. Besides,  (Wang et al., 2022; Zhang et al., 2023) leverage the invariant learning to address popularity debiasing, and  (Wang et al., 2024) captures the invariant KG subgraph to enhance recommendations. In this work, we introduce the invariant learning principle to the social recommendation, intending to capture stable user preferences that persist across diverse social environments.
@@ 4.2. Invariant Preference Learning
After obtaining the learned node representations in the $k^{th}$ social environment, we can infer the propensity that user $a$ will interacts with item $i$ by an inner product: $\hat{r}_{ai}=<p_{a}^{k},q_{i}^{k}>$. All the above process are summarized as $\hat{r}_{ai}^{k}=f_{\phi}^{k}(a,i,\mathcal{G}_{r},\mathcal{G}_{s}^{k})$, where $\phi=\{\mathbf{P}^{0},\mathbf{Q}^{0}\}$ denote the initialized embedding parameters.
@@
4.3. Model Discussion
In this section, we analyze the proposed SGIL from both space and time complexity.
4.3.1. Space Complexity
As illustrated in Algorithm 1, the parameters of SGIL consist of two components: the embedding parameters $\phi$ and the environment generator parameters $\theta$. The embedding parameters $\phi\in\mathbb{R}^{(M+N)\times d}$ are the same as those used in general social recommenders (such as DiffNet(Wu et al., 2019) and GBSR (Yang et al., 2024)). In addition, the environment generator parameters $\theta$ consist of $K$ individual MLPs ($\mathbb{R}^{2d\times d}+\mathbb{R}^{d\times 1}$). Since the embedding size $d$ is much smaller than node number $M+N$, the additional parameters introduced by SGIL is negligible.
4.3.2. Time Complexity
The computational overhead introduced by SGIL mainly consists of three components: environment simulation, invariant preference learning, and adversarial training. For environment simulation, we first obtain the encoded node representations using GCNs, and then employ $K$ individual MLPs to generate diverse environments. For preference learning, we learn user and item representations within each environment and optimize them through an invariance-based objective. Finally, we adopt an across-batch adversarial training strategy to iteratively update the parameters. Compared to the backbone model (LightGCN-S), SGIL incurs approximately $K$ times the computational cost during the preference learning stage, and the employed softmax-base loss(Wu et al., 2024) is computationally more intensive than the traditional BPR loss (Rendle et al., 2009). Fortunately, in our experiments, the optimal value of $K$ is typically less than 5, and SGIL converges much faster than LightGCN-S. Overall, the total training time of SGIL is approximately 1 to 3 times that of LightGCN-S.
@@ 5.1.1. Datasets
The evaluation metrics are Recall@N and NDCG@N (Järvelin and Kekäläinen, 2002), which have been widely used in ranking-based recommender systems. Following the mainstream evaluation process (Zhao et al., 2020), we adopt a full-ranking protocol that all non-interacted items are candidates for evaluation.
@@ 5.1.3. Implementation Details
All embedding-based methods are initialized by a normal initializer with a mean value of 0 and a standard variance of 0.01, the embedding size is fixed to 64. For model optimization, we use the Adam optimizer (Kingma and Ba, 2014) with a learning rate of 0.001 and a batch size of 2048. We empirically set the GCN layer $L=3$ and the temperature coefficient $\tau=0.2$ for all datasets. For the edge observation bias, we follow the previous work (Yang et al., 2024) and set the bias $\epsilon=0.5$ for all datasets. For the environment number $K$, the variance-based penalty coefficient $\beta$, we carefully search the best parameters for each dataset and report the detailed comparisons. For the adversarial training step, we set $T=20$ on both the Douban-Book and Yelp datasets, and $T=3$ on the Epinions datasets. For all baselines, we refer to the original parameters and carefully fine-tune them for fair comparisons.
@@
5.2. Performance Comparisons
Following the current SOTA method (Yang et al., 2024), we implement all social denoising methods on the LightGCN-S backbone. As shown in Table 2, we have the following observations:
@@ 5.3. Ablation Study
5.4. Detailed Analysis of SGIL
@@ 6. Conclusion
Acknowledgements.
This work was supported in part by grants from the National Key Research and Development Program of China (Grant No. 2021ZD0111802), the National Natural Science Foundation of China (Grant No. U23B2031, 721881011).